HETAV PATEL

Grade 11,

. 25kg : 20g = 50kg : 40g . 25kg : 20g = 50kg : 40g

. 25kg : 20g = 50kg : 40g
. 25kg : 20g = 50kg : 40g

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The given statement is True.
We know that, 1 kg = 1000 g So, 25 × 1000 = 25000 g

50 × 1000 = 50000g
Then, 25000g : 20g = 50000g : 40g

25000/20 = 50000/40

2500/2 = 5000/4
2500/2 is further simplified by dividing both numerator and denominator by 2 we get, = 1250
5000/4 is further simplified by dividing both numerator and denominator by 4 we get, = 1250
Hence, 1250 = 1250
So, 25kg : 20g = 50kg : 40g


Thanks